# MULTIBENCH++: A Unified and Comprehensive Multimodal Fusion Benchmarking Across Specialized Domains

### v1

@@ 1 Introduction
Affective Computing MELD $\{a,t\}$ 13,708 Emotion/Sentiment Recognition
&Social Media IEMOCAP $\{a,v,t\}$ 7,433 Emotion/Sentiment Recognition
Understanding MAMI $\{i,t\}$ 11,000 Misogyny Content Detection
@@
The field of multimodal fusion is growing rapidly, but its progress is held back by a key bottleneck: the lack of a comprehensive and standard evaluation system. Currently, the vast majority of new models still depend on a small number of classic datasets for their development and validation. However, in terms of their scale, diversity, and complexity, these datasets fall short of representing the dynamic challenges found in the real world. This reliance causes two main problems: models tend to overfit to the biases of these specific datasets, which limits their generalization to new scenarios, and it makes fair, objective comparisons between different fusion methods very difficult. Four years ago, the groundbreaking MULTIBENCH framework partially addressed this issue by providing a unified, cross-domain evaluation platform. Yet, with the fast pace of change in the field, the existing MULTIBENCH is no longer sufficient to meet today’s higher standards for benchmarks, and its limitations are becoming more obvious.
The urgent need to create a new generation of evaluation benchmarks comes from two main trends in the multimodal field. The first trend is the explosive growth in the amount and complexity of data. With the widespread use of internet content, IoT devices, and applications in specialized fields like medical imaging and autonomous driving (Hu et al. 2023; Kong et al. 2011), multimodal data is now much larger, more diverse, and often “noisier” or less perfect. This fundamental shift means that older benchmarks, can no longer show how robust a model is or how well it can adapt to complex, real-world situations. The second trend is that the models for fusing data are changing quickly and becoming more varied, especially with the rise of Transformer-based methods. Driven by advances in deep learning, researchers have introduced a huge variety of new fusion models (Wei et al. 2020; Wang et al. 2022; Xu, Zhu, and Clifton 2023). However, without a standard way to test them, these models are often evaluated on different datasets, making it nearly impossible to fairly compare which methods truly work best. Therefore, building a standardized platform that can thoroughly and fairly test different models against real-world challenges has become essential.

### v2

@@ 1 Introduction
Affective Computing &Social Media Understanding MELD $\{a,t\}$ 13,708 Emotion/Sentiment Recognition
IEMOCAP $\{a,v,t\}$ 7,433 Emotion/Sentiment Recognition
MAMI $\{i,t\}$ 11,000 Misogyny Content Detection
@@
However, a significant divergence exists between multimodal research and other domains. While fields such as natural language processing and graph learning have successfully converged on dominant architectural paradigms, specifically the Transformer and GNNs, the multimodal domain conspicuously lacks an equivalent unified, foundational framework. Progress remains highly fragmented; researchers typically validate new methods on a small, bespoke selection of classic datasets. This reliance on siloed benchmarks is a key bottleneck. It not only leads to models overfitting to specific data biases and prevents fair objective comparisons, but more importantly, it has hindered the systematic search for a truly general-purpose fusion architecture. Four years ago, the groundbreaking MULTIBENCH framework partially addressed this by providing a unified evaluation platform. Yet, with the field’s rapid evolution, its limitations are now apparent, and it is no longer sufficient to meet today’s challenges.
The urgent need for a new foundational platform stems from two primary trends. The first is the explosive growth in data combinatorial complexity. Unlike unimodal tasks, real-world applications in medical imaging, IoT, and autonomous driving (Hu et al. 2023; Kong et al. 2011) span a vast, heterogeneous spectrum. This combinatorial effect, where different data combinations can yield entirely different analytical conclusions, means older, simpler benchmarks can no longer demonstrate a model’s robustness or adaptability to real-world complexity. The second trend is the rapid evolution of fusion models, especially Transformer-based methods (Wei et al. 2020; Wang et al. 2022; Xu, Zhu, and Clifton 2023). Without a standardized, complex testbed, it is impossible to determine if these new models are truly general-purpose or simply adept at specific data combinations. Therefore, a platform that forces models to confront this combinatorial complexity has become essential to guide the search for a foundational architecture.